find two consecutive odd integers such that six times the smaller is three more than five times the larger
Answer by JulietG(1812)

Here's what we know:
A = B+2 (consecutive odd integers), where B is the smaller number
6B - 3 = 5A (six times the smaller is three more than five times the larger )
Substitute the value of A from the first equation into the second.
6B-3 = 5(B+2)
Distribute
6B-3 = 5B+10
Add 3 to each side
6B = 5B + 13
Subtract 5B from each side
B = 13
.
If B is 13, then A is 15. Let's prove it.
6*13 = (5*15)+3
78 = 75+3
78=78
Success~!